OpenStudy (anonymous):

solve the system using elimination3x - 2y = 0 4x + 2y = 14

OpenStudy (anonymous):

Eliminate either the x or y term so that you are left with on variable equaling a number. You eliminate the variables by subtracting the equations with each other.

OpenStudy (shaik0124):

7x=14 x=14/2 x=7 now from 3*7-2y=0 21-2y=0 21=2y y=21/2

OpenStudy (shaik0124):

@lovechama got it

OpenStudy (shaik0124):

just u add the two equations first then u will get x and substitute x in anyone equation u will get y

OpenStudy (anonymous):

how u get x=7 if u divide 7 in the 14 whish equal 2

OpenStudy (shaik0124):

sorry u will x=14/7 x=2 in hurry it was typing mistake

OpenStudy (anonymous):

it okay

OpenStudy (anonymous):

dont y suppose equal 7 cuz 3 divide in 7 is 21